# NODEC: Neural ODE For Optimal Control of Unknown Dynamical Systems

### 03 Jan 2024 v1

@@ Abstract
Controlling complex dynamical systems is generally associated with minimizing certain control objectives with known dynamics under the variational calculus framework. For systems with unknown dynamics, an additional step of dynamics modeling is required. However, any inaccuracy in dynamics modeling will lead to sub-optimality in the resulting control function. Another set of approaches for controlling unknown dynamical systems - reinforcement learning, folds the dynamics modeling into controller training via value function approximation or policy gradient through extensively interacting with the environment, but it suffers from low data efficiency. To address these, we introduce NODEC, a novel framework for controlling unknown dynamical systems, which combines dynamics modelling and controller training using a coupled neural ODE model. Through an intriguing interplay between the two coupled neural networks, NODEC learns system dynamics as well as optimal controls that guides the unknown dynamical system towards target states. Our experiments demonstrate the effectiveness and data efficiency of NODEC for learning optimal control of unknown dynamical systems.
Keywords Neural ODE, Dynamical System, Optimal Control
@@ 1 Introduction
Optimal control is a robust mathematical framework that enables finding control and state trajectories for dynamical systems over a period of time to minimize a performance metric [1]. The primary objective in optimal control is to determine a control law that minimizes or maximizes a performance metric, such as energy consumption, time to reach a target, or cost of operation. Solving optimal control problems can be difficult, especially for systems with nonlinear dynamics. Various methods have been developed to solve optimal control problems, including principle-driven approaches such as Pontryagin’s maximum principle [2] [3], and data-driven approach based on modern machine learning and optimization techniques [4] [5] [6].
Our approach discussed in this paper falls into the data-driven category. In particular, we introduce NODEC, a novel control framework based on coupled neural ordinary differential equations that automatically learns optimal control and system dynamics for unknown dynamical systems.
1.1 Principle driven optimal control
@@
In this section, we briefly explain the principle-driven approach for solving optimal control problems and particularly the Maximum principle [2].
@@
where $\lambda^{*}(t)$ are the costate variables associated with $x^{*}(t)$. The aforementioned optimality condition typically leads to a system of differentiable equations, and the solutions to this system contain the desired optimal control. This principle-driven approach is often analytically and computationally intractable when applied to complex dynamical systems. With the recent development in differentiable programming and physics-informed neural network [7], the maximum principle based approaches can be extended beyond analytical solutions and include higher-order terms in the optimality condition [8].
1.2 Data driven optimal control
Another class of optimal control methods relies more on differentiable optimization and less on theoretical optimality conditions. For instance, reinforcement learning is an algorithmic framework for solving optimal control problems - generally in an unknown environment. Even though certain RL algorithms can be used to approximate the solution to the Hamilton-Jacobi-Bellman (HJB) equation (necessary and sufficient condition for optimality), recent advancements in RL such as the policy gradient-based methods [9] [10] or the actor-critic based methods [11] are in a purely data-driven regime. They directly parameterize the control policy and collect a large number of interactions with the environment to build differentiable loss for improving the parameterized policy, thus skipping the dynamics modelling and the derivation of optimality condition as in the principle-driven approach. A subroutine in RL that involves dynamics modelling is known as the model-based RL, where the approximation of unknown dynamics amounts to a supervised learning problem, and the learned dynamics act as a subroutine in the original RL policy optimization process in the form of forward planning [12]. However, powerful policy optimizers exploit inaccuracies in learned dynamical models leading to adversarial examples and sub-optimal policies [13]. RL algorithms are known for having a low data efficiency [14], and the available data can be extremely limited for some dynamical systems [15].
A closely related work is done by [6], where they embed a parameterized controller inside the real dynamical functions $f$, and show that it is possible to generate control signals that resemble those of optimal control by minimizing a loss without relying on solving maximum principle or HJB equations. However, their work is strictly limited to controlling dynamical systems with known dynamical functions.
2 Method
In this section, we introduce our method NODEC: Neural ODE Controller, a coupled neural ODE structure that simultaneously learns optimal controls as well as dynamics of an unknown dynamical system.
@@ 2.1 Neural ODE
2.2 Architecture of NODEC
NODEC consists of a coupled neural ODE structure. We still parameterize the control function $u(x,t)$ through a neural network $h_{\theta}(x,t)$, denoted as the controller. This controller is embedded as part of the dynamics represented by another neural network $g_{\gamma}(x,u,t)$ denoted as the dynamics learner, and the control signal $u$ is given by the controller. This is represented by the following equations:
@@ 2.2 Architecture of NODEC
This is essentially a coupled neural ODE where the neural part consists of two neural networks: controller $h_{\theta}$ and dynamics learner $g_{\gamma}$. The integration in Equation (5) is conducted using a numerical solver over the dynamics specified by the coupled $h_{\theta}$ and $g_{\gamma}$, and this is how NODEC makes a prediction about the state $x_{T}$ as a result of continuously controlling the state evolution using the controller in the environment with unknown dynamics. For instance, starting at $x_{0}$, NODEC predicts where the system will end up by:
@@
2.3 Training Loss
There are two objectives when we train our NODEC model given by ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. The first objective is for the dynamics learner $\hat{g}_{\gamma}$ to approximate the real unknown dynamics $f(x,u,t)$: the integration in Equation (5) gives us the model prediction of the trajectory roll-out, and we want this prediction to match the real trajectory coming from the environment (the real dynamics) using the same controller ${h}_{\theta}$. The second objective is for the controller ${h}_{\theta}$: we want to satisfy the optimal control objectives, for example, by continuously applying control signal u given by ${h}_{\theta}$, the end state in the predicted trajectory $x_{T}$ using the dynamics learner ${g}_{\gamma}$ (which is also the end state in the real trajectory if ${g}_{\gamma}$ is well trained) should be close to our target state $x^{*}$.
@@ 2.3 Training Loss
2.4 Alternative Training
A naive approach for training NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$ is a two-stage training process. In the first stage, the dynamics learner ${g}_{\gamma}$ is trained to accurately capture the underlying dynamics $f$ by leveraging a diverse set of controllers (ideally all possible control functions) for generation the input controls u. In the second stage, the controller ${h}_{\theta}$ is trained within the learned dynamics model ${g}_{\gamma}$ to attain optimal control, which is similar to the approach described in [6] for learning optimal control with known dynamics.
@@ 2.4 Alternative Training
This naturally leads to the alternative training scheme depicted in Figure (2) for training NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. In this scheme, we alternate the training of the two neural networks, ${g}_{\gamma}$ and ${h}_{\theta}$, for in total of K iterations. We denote the neural networks at the $k_{th}$ alternation as ${g}_{\gamma}^{k}$ and ${h}_{\theta}^{k}$, respectively.
@@
Algorithm 1 Alternative Training of NODEC Model
@@ 3.1 Control task: Linear System
We now investigate the control performance of NODEC applied to a linear dynamical system described by $f(x,u)=Ax+Bu$.
@@
Figure 3: The full training process of NODEC for Ax+Bu task
Figure 3: The full training process of NODEC for Ax+Bu task
@@
This control task is the same task [6] used, but we assume the dynamics is unknown, and our NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$ does not have access to the true dynamics $f(x,u)$. We aim to evaluate whether our NODEC can learn the analytical optimal control without the knowledge of the true dynamics. We set problem dynamics parameters: A = $\begin{pmatrix}01\\ 10\end{pmatrix}$ and the driver matrix B = $\begin{pmatrix}1\\ 0\end{pmatrix}$. We use a batch of uniformly random five initial states, the target state is set to be $x^{*}=(1,-1)$, and control horizon is set to be $T=1$. Therefore, our control objective is to bring the system state $x_{T}$ at $T=1$ closer to the target state starting from those initial states $x_{0}$ in the randomly generated batch.
@@
For solving this linear control task with our NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$, we employ a controller ${h}_{\theta}$ and a dynamics learner ${g}_{\gamma}$ with one hidden layer consisting of 30 neurons each. The output layer of both neural networks uses the identity activation function, while the hidden layer uses the tanh non-linear activation function.
@@ Results
Figure (3) demonstrate the full training process of NODEC when controlling the $Ax+Bu$ linear system. It provides an overview of the full alternative training process, consisting of six pairs of plots separated by dashed lines. Each pair corresponds to the results obtained in a specific training alternation $k$. The top-left pair corresponds to the initial training alternation ($k=0$), while the bottom-right pair represents the fully trained model after $K=10$ alternations between the dynamics learner and the controller training. The pairs in between correspond to intermediate training alternations.
@@
The final trajectory plot at the bottom right corner reveals that the trained NODEC controller $h_{\theta}^{K}$ successfully guides the system from various initial states to the target state $x^{*}=(1,-1)$, in both learned dynamics and real dynamics (which shows that the training of the $g_{\gamma}^{K}$ is successful as well). The training losses for dynamics learner and controller corresponds to these six alternations (each alternation we take $n_{c}$, $n_{g}$ gradient steps) are shown in Figure (8) Appendix C.
Moreover, it is fascinating to observe that the controller and dynamics learner converge almost simultaneously: as the controller performs better (end states in trajectories are closer to the target state), the dynamics learner adjusts to fit the dynamics more accurately (deviation between black and red trajectories are smaller). As discussed in Section 2.4, an intriguing interaction unfolds between the controller and the dynamics learner. The controller effectively guides the dynamics learner’s learning process by leading it to different regions of the vector field in $f$ that require accurate fitting. Conversely, the dynamics learner imposes constraints on the controller by specifying the dynamics that the controller will operate within. In essence, they mutually shape each other’s training objectives and supervise one another during each alternation $k$. As the alternations progress, the quality of this mutual supervision between the two neural networks continually improves. As the controller’s performance enhances, the dynamics learner learns to approximate dynamics that are closer to optimal trajectories in the state space. Simultaneously, as the dynamics learner becomes more proficient, the controller’s training becomes more effective in terms of the objective of steering the states closer to the target states, as the dynamics that the controller operates in is closer to the true dynamics. This iterative feedback loop reinforces their collaborative training and leads to increasingly refined coordination between the controller and the dynamics learner.
These findings once again highlight the captivating interactions and relationships between the controller $h_{\theta}$ and dynamics learner $g_{\gamma}$ in our NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. Through this enthralling interplay between the two neural networks, we can solve optimal control tasks without any prior knowledge of the system dynamics.
@@ 3.2 Control Task: CartPole
We now investigate the control performance of NODEC applied to a non-linear dynamical system known as CartPole or inverted pendulum.
@@
Figure 5: The full training process of NODEC for CartPole task
Figure 5: The full training process of NODEC for CartPole task
An important observation is that the control loss in our NODEC model for the CartPole task has transitioned from a single-end-state loss to an integral loss. The control loss now measures the mean squared error between the desired state (0,0,0,0) and all states along the trajectories. This is because the objective is to maintain the pole in an upright position throughout the entire trajectory. This can be easily incorporated into our NODEC alternative training algorithm by changing line 17 of Algorithm 1.
@@
For solving the CartPole control task with our NODEC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$, we set up our controller ${h}_{\theta}$ with a single hidden layer consists of 64 neurons and one-dimensional output for control F, and our dynamics learner ${g}_{\gamma}$ with two hidden layers consisting of 64 and 32 neurons respectively and having a four-dimensional output corresponds to $\dot{x},\ddot{x},\dot{\theta},\ddot{\theta}$. The output layer for both ${h}_{\theta}$ and ${g}_{\gamma}$ uses the identity activation function, while the hidden layer uses the tanh non-linear activation function. We set the training alternation parameter to $K=12$ in Algorithm 1. During each alternation, we train the controller for $n_{c}=2000$ gradient steps with a learning rate of 0.0001. The training of the dynamic learner follows a multi-stage approach that enhances the stability of training. Firstly, we minimize the mean squared errors of trajectory roll-outs from $t=0$ to $t=0.5$ for 2000 gradient steps, using a learning rate of 0.005. Next, we extend the training to cover the time range from $t=0$ to $t=1$ for an additional 2000 gradient steps, with a learning rate of 0.001. Finally, we continue the training from $t=0$ to $t=T$ for another 2000 gradient steps, employing a learning rate of 0.0001. This multi-stage training can be seen in Figure (9) Appendix C, where the dynamics learner loss plots have spikes for every 2000 gradient steps. On the other hand, the dynamic trainer follows a multi-stage training approach for the purpose of stability. AT first, the time step is set to $t=0.5$ with $n_{c}=10000$ gradient steps and a learning rate of 0.001. Then, the time step is extended to a time range of $t=0.5$ to $t=1$ with an additional 10000 gradient steps $n_{c}$ and a learning rate of 0.0001.

### 04 Feb 2024 v3

# Neural Control: Concurrent system identification and control learning with neural ODE

@@ Abstract
Controlling continuous-time dynamical systems is generally a two step process: first, identify or model the system dynamics with differential equations, then, minimize the control objectives to achieve optimal control function and optimal state trajectories. However, any inaccuracy in dynamics modeling will lead to sub-optimality in the resulting control function. To address this, we propose a neural ODE based method for controlling unknown dynamical systems, denoted as Neural Control (NC), which combines dynamics identification and optimal control learning using a coupled neural ODE. Through an intriguing interplay between the two neural networks in coupled neural ODE structure, our model concurrently learns system dynamics as well as optimal controls that guides towards target states. Our experiments demonstrate the effectiveness of our model for learning optimal control of unknown dynamical systems.
Keywords Neural ODE, Dynamical System, Optimal Control, System identification
@@ 1 Introduction
Optimal control is a robust mathematical framework that enables finding control and state trajectories for dynamical systems over a period of time to minimize a performance metric [1]. The primary objective in optimal control is to determine a control law that minimizes or maximizes a performance metric, such as energy consumption, time to reach a target, or cost of operation. Solving optimal control problems can be difficult, especially for systems with nonlinear dynamics. Classical optimal control theory, a cornerstone of modern control systems engineering, has evolved through seminal contributions that establish principles for optimizing dynamic system behaviors. Central to this evolution is the calculus of variations [2], providing foundational techniques for formulating and solving optimization problems in continuous systems. Another pivotal contribution is Pontryagin’s Maximum Principle [3] [4] offering necessary conditions for optimality in control problems, widely applicable in fields from aerospace to economics. The Hamilton-Jacobi-Bellman (HJB) equation, foundational in dynamic programming, extends these concepts to more complex systems [5]. These works collectively underpin the theoretical framework of classical optimal control.
While classical approaches to optimal control are powerful, they often hinge on the analytical resolution of complex equations, a task that can be daunting and computationally intensive. Recent advancements have seen the integration of Machine Learning methods, particularly artificial neural networks (ANNs), into the domain of optimal control. Works such as those by authors in [6] and [7] demonstrate the potential of embedding parameterized control functions within a system’s dynamical equations. These approaches leverage auto-differentiation capabilities of ANNs to iteratively refine and optimize control functions, a significant leap from traditional methods.
However, a critical assumption underlying these AI-enhanced methods is the prior knowledge of system dynamics. Typically, this knowledge is encapsulated in differential equations, formulated through system identification – a process of building mathematical models from empirical data. Traditionally, system identification relies on statistical methods as elucidated in [8]. The advent of modern machine learning techniques has also revolutionized this process, enhancing both efficiency and accuracy in modeling dynamic systems. Pioneering works in this area, such as those by [9], [10], and [11], exemplify the successful application of these advanced techniques in system identification tasks.
Building on these developments, we introduce - Neural Control (NC), which synergizes system identification with the learning of optimal control, specifically tailored for continuous time dynamical systems. This framework eliminates the need for separate system identification prior to control optimization in optimal control framework. By employing a coupled neural ordinary differential equations, NC learns the underlying system dynamics while simultaneously learning the optimal control function for continuous time optimal control.
Progress in automatic differentiation and machine learning methodologies has significantly enhanced the evolution of modeling and control strategies in the domain of optimal control. For instance, reinforcement learning is an algorithmic framework that can be applied for optimal control problems - generally in an unknown environment. Even though certain RL algorithms can be used to approximate the solution to the Hamilton-Jacobi-Bellman (HJB) equation (necessary and sufficient condition for optimality), recent advancements in RL such as the policy gradient-based methods [12] [13] or the actor-critic based methods [14] are in a purely data-driven regime. They directly parameterize the control policy and collect a large number of interactions with the environment to build differentiable loss for improving the parameterized policy, thus skipping the dynamics modelling and the derivation of optimality condition as in the principle-driven approach. A subroutine in RL that involves dynamics modelling is known as the model-based RL, where the approximation of unknown dynamics amounts to a supervised learning problem, and the learned dynamics act as a subroutine in the original RL policy optimization process in the form of forward planning [15]. However, as dynamics modeling and policy learning are often separate processes, powerful policy optimizers exploit inaccuracies in learned dynamical models leading to adversarial examples and sub-optimal policies [16]. Besides, RL algorithms are known for having a low data efficiency [17], and the available data can be extremely limited for some dynamical systems [18].
Neural Ordinary Differential Equations (Neural ODEs) have demonstrated considerable efficacy as a framework for modeling continuous systems using time series data that is sampled irregularly [19]. There has been a growing interest in the application of Neural ODEs within the realm of continuous-time control tasks. Notably, the use of Neural ODEs has been explored in the domain of learning optimal control functions, as evidenced in studies such as [6], [7], [20], and [21]. These investigations have incorporated a parameterized controller within the actual dynamical functions $f$, revealing that such controllers can emulate optimal control signals by minimizing a loss function, circumventing the explicit resolution of maximum principle or Hamilton-Jacobi-Bellman (HJB) equations.
Despite these advancements, existing methodologies predominantly focus on controlling continuous-time dynamical systems when the dynamical functions are known. Conversely, the inherent suitability of Neural ODEs for modeling continuous dynamical systems has prompted their extensive application in learning the dynamics of the underlying system or system identification tasks, as highlighted in [9], [22], and [23].
However, to the best of our knowledge, there has been no methodological integration of system identification and the learning of optimal control policies within a singular Neural ODE framework. In response to this gap, we introduce Neural Control (NC). NC represents a novel approach that combines the dual processes of learning the optimal control function and identifying system dynamics. This integration is executed within a coupled Neural ODE structure, enabling concurrent optimal control function learning and system identification in an end-to-end manner. This approach aims to harness the full potential of the Neural ODE model for controlling continuous-time systems.
2 Method
In this section, we introduce our method: Neural Control (NC), a coupled neural ODE structure that simultaneously learns optimal control as well as system dynamics of unknown dynamical systems. We first briefly introduce few relevant concepts before discussions of NC.
2.1 Classical optimal control
@@
In this section, we briefly explain the classical methods for solving optimal control problems and particularly the Maximum principle [3].
@@
where $\lambda^{*}(t)$ are the costate variables associated with $x^{*}(t)$. The aforementioned optimality condition typically leads to a system of differentiable equations, and the solutions to this system contain the desired optimal control. This principle-driven approach is often analytically and computationally intractable when applied to complex dynamical systems.
@@ 2.2 Neural ODE
2.3 Neural Control: NC
2.3.1 Model architecture
Our NC model consists of a coupled neural ODE structure. We still parameterize the control function $u(x,t)$ through a neural network $h_{\theta}(x,t)$, denoted as the controller. This controller is then embedded as part of the dynamics represented by another neural network $g_{\gamma}(x,u,t)$ denoted as the dynamics learner, and the control signal $u$ is given by the controller. This is represented by the following equations manifested by a coupled neural ODE structure:
@@ 2.3.1 Model architecture
This is essentially a coupled neural ODE where the neural part consists of two neural networks: controller $h_{\theta}$ and dynamics learner $g_{\gamma}$. The integration in Equation (5) is conducted using a numerical solver over the dynamics specified by the coupled $h_{\theta}$ and $g_{\gamma}$, and this is how NC makes a prediction about the state $x_{T}$ as a result of continuously controlling the state evolution using the controller in the environment with unknown dynamics. For instance, starting at $x_{0}$, NC predicts where the system will end up by:
@@
2.3.2 Training Loss
There are two objectives when we train our NC model given by ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. The first objective is for the dynamics learner $\hat{g}_{\gamma}$ to approximate the real unknown dynamics $f(x,u,t)$: the integration in Equation (5) gives us the model prediction of the trajectory roll-out, and we want this prediction to match the real trajectory coming from the environment (the real dynamics) using the same controller ${h}_{\theta}$. The second objective is for the controller ${h}_{\theta}$: we want to satisfy the optimal control objectives, for example, by continuously applying control signal u given by ${h}_{\theta}$, the end state in the predicted trajectory $x_{T}$ using the dynamics learner ${g}_{\gamma}$ (which is also the end state in the real trajectory if ${g}_{\gamma}$ is well trained) should be close to our target state $x^{*}$.
@@ 2.3.2 Training Loss
2.3.3 Alternative Training
A naive approach for training NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$ is a two-stage training process. In the first stage, the dynamics learner ${g}_{\gamma}$ is trained to accurately capture the underlying dynamics $f$ by leveraging a diverse set of controllers (ideally all possible control functions) for generation the input controls u. In the second stage, the controller ${h}_{\theta}$ is trained within the learned dynamics model ${g}_{\gamma}$ to attain optimal control, which is similar to the approach described in [6] for learning optimal control with known dynamics.
@@ 2.3.3 Alternative Training
This naturally leads to the alternative training scheme depicted in Figure (2) for training NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. In this scheme, we alternate the training of the two neural networks, ${g}_{\gamma}$ and ${h}_{\theta}$, for in total of K iterations. We denote the neural networks at the $k_{th}$ alternation as ${g}_{\gamma}^{k}$ and ${h}_{\theta}^{k}$, respectively.
@@
Algorithm 1 Alternative Training of NC Model
@@ 3.1 Control task: Linear System
We now investigate the control performance of NC applied to a linear dynamical system described by $f(x,u)=Ax+Bu$. This is the same tasked used in [6] and [26].
@@
Figure 3: The full training process of NC for Ax+Bu task
Figure 3: The full training process of NC for Ax+Bu task
@@
This control task is the same task [6] used, but we assume the dynamics is unknown, and our NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$ does not have access to the true dynamics $f(x,u)$. We aim to evaluate whether our NC can learn the analytical optimal control without the knowledge of the true dynamics. We set problem dynamics parameters: A = $\begin{pmatrix}01\\ 10\end{pmatrix}$ and the driver matrix B = $\begin{pmatrix}1\\ 0\end{pmatrix}$. We use a batch of uniformly random five initial states, the target state is set to be $x^{*}=(1,-1)$, and control horizon is set to be $T=1$. Therefore, our control objective is to bring the system state $x_{T}$ at $T=1$ closer to the target state starting from those initial states $x_{0}$ in the randomly generated batch.
@@
For solving this linear control task with our NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$, we employ a controller ${h}_{\theta}$ and a dynamics learner ${g}_{\gamma}$ with one hidden layer consisting of 30 neurons each. The output layer of both neural networks uses the identity activation function, while the hidden layer uses the tanh non-linear activation function.
@@ Results
Figure (3) demonstrate the full training process of NC when controlling the $Ax+Bu$ linear system. It provides an overview of the full alternative training process, consisting of six pairs of plots separated by dashed lines. Each pair corresponds to the results obtained in a specific training alternation $k$. The top-left pair corresponds to the initial training alternation ($k=0$), while the bottom-right pair represents the fully trained model after $K=10$ alternations between the dynamics learner and the controller training. The pairs in between correspond to intermediate training alternations.
@@
The final trajectory plot at the bottom right corner reveals that the trained NC controller $h_{\theta}^{K}$ successfully guides the system from various initial states to the target state $x^{*}=(1,-1)$, in both learned dynamics and real dynamics (which shows that the training of the $g_{\gamma}^{K}$ is successful as well). The training losses for dynamics learner and controller corresponds to these six alternations (each alternation we take $n_{c}$, $n_{g}$ gradient steps) are shown in Figure (8) Appendix C.
Moreover, it is fascinating to observe that the controller and dynamics learner converge almost simultaneously: as the controller performs better (end states in trajectories are closer to the target state), the dynamics learner adjusts to fit the dynamics more accurately (deviation between black and red trajectories are smaller). As discussed in Section 2.3.3, an intriguing interaction unfolds between the controller and the dynamics learner. The controller effectively guides the dynamics learner’s learning process by leading it to different regions of the vector field in $f$ that require accurate fitting. Conversely, the dynamics learner imposes constraints on the controller by specifying the dynamics that the controller will operate within. In essence, they mutually shape each other’s training objectives and supervise one another during each alternation $k$. As the alternations progress, the quality of this mutual supervision between the two neural networks continually improves. As the controller’s performance enhances, the dynamics learner learns to approximate dynamics that are closer to optimal trajectories in the state space. Simultaneously, as the dynamics learner becomes more proficient, the controller’s training becomes more effective in terms of the objective of steering the states closer to the target states, as the dynamics that the controller operates in is closer to the true dynamics. This iterative feedback loop reinforces their collaborative training and leads to increasingly refined coordination between the controller and the dynamics learner.
These findings once again highlight the captivating interactions and relationships between the controller $h_{\theta}$ and dynamics learner $g_{\gamma}$ in our NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$. Through this enthralling interplay between the two neural networks, we can solve optimal control tasks without any prior knowledge of the system dynamics.
@@ 3.2 Control Task: CartPole
We now investigate the control performance of NC applied to a non-linear dynamical system known as CartPole or inverted pendulum.
@@
Figure 5: The full training process of NC for CartPole task
Figure 5: The full training process of NC for CartPole task
An important observation is that the control loss in our NC model for the CartPole task has transitioned from a single-end-state loss to an integral loss. The control loss now measures the mean squared error between the desired state (0,0,0,0) and all states along the trajectories. This is because the objective is to maintain the pole in an upright position throughout the entire trajectory. This can be easily incorporated into our NC alternative training algorithm by changing line 17 of Algorithm 1.
@@
For solving the CartPole control task with our NC model ${g}_{\gamma}(x,{h}_{\theta}(x,t),t)$, we set up our controller ${h}_{\theta}$ with a single hidden layer consists of 64 neurons and one-dimensional output for control F, and our dynamics learner ${g}_{\gamma}$ with two hidden layers consisting of 64 and 32 neurons respectively and having a four-dimensional output corresponds to $\dot{x},\ddot{x},\dot{\theta},\ddot{\theta}$. The output layer for both ${h}_{\theta}$ and ${g}_{\gamma}$ uses the identity activation function, while the hidden layer uses the tanh non-linear activation function. We set the training alternation parameter to $K=12$ in Algorithm 1. During each alternation, we train the controller for $n_{c}=2000$ gradient steps with a learning rate of 0.0001. The training of the dynamic learner follows a multi-stage approach that enhances the stability of training. Firstly, we minimize the mean squared errors of trajectory roll-outs from $t=0$ to $t=0.5$ for 2000 gradient steps, using a learning rate of 0.005. Next, we extend the training to cover the time range from $t=0$ to $t=1$ for an additional 2000 gradient steps, with a learning rate of 0.001. Finally, we continue the training from $t=0$ to $t=T$ for another 2000 gradient steps, employing a learning rate of 0.0001. This multi-stage training can be seen in Figure (9) Appendix C, where the dynamics learner loss plots have spikes for every 2000 gradient steps. On the other hand, the dynamic trainer follows a multi-stage training approach for the purpose of stability. AT first, the time step is set to $t=0.5$ with $n_{c}=10000$ gradient steps and a learning rate of 0.001. Then, the time step is extended to a time range of $t=0.5$ to $t=1$ with an additional 10000 gradient steps $n_{c}$ and a learning rate of 0.0001.